# A Sanity Check for AI-generated Image Detection

### v2

@@ Abstract
With the rapid development of generative models, discerning AI-generated content has evoked increasing attention from both industry and academia. In this paper, we conduct a sanity check on “whether the task of AI-generated image detection has been solved”. To start with, we present Chameleon dataset, consisting of AI-generated images that are genuinely challenging for human perception. To quantify the generalization of existing methods, we evaluate 9 off-the-shelf AI-generated image detectors on Chameleon dataset. Upon analysis, almost all models misclassify AI-generated images as real ones. Later, we propose AIDE (AI-generated Image DEtector with Hybrid Features), which leverages multiple experts to simultaneously extract visual artifacts and noise patterns. Specifically, to capture the high-level semantics, we utilize CLIP to compute the visual embedding. This effectively enables the model to discern AI-generated images based on semantics and contextual information; Secondly, we select the highest and lowest frequency patches in the image, and compute the low-level patchwise features, aiming to detect AI-generated images by low-level artifacts, for example, noise patterns, anti-aliasing, etc. While evaluating on existing benchmarks, for example, AIGCDetectBenchmark and GenImage, AIDE achieves +3.5% and +4.6% improvements to state-of-the-art methods, and on our proposed challenging Chameleon benchmarks, it also achieves the promising results, despite this problem for detecting AI-generated images is far from being solved.
@@ 1 Introduction
Recently, the vision community has witnessed remarkable advancements in generative models. These methods, ranging from generative adversarial networks (GANs) (Goodfellow et al., 2014; Zhu et al., 2017; Brock et al., 2018; Karras et al., 2019) to diffusion models (DMs) (Ho et al., 2020; Nichol & Dhariwal, 2021; Rombach et al., 2022; Song et al., 2020; Liu et al., 2022b; Lu et al., 2022; Hertz et al., 2022; Nichol et al., 2021) have demonstrated unprecedented capabilities in synthesizing high-quality images that closely resemble real-world scenes. On the positive side, such generative models have enabled various valuable tools for artists and designers, democratizing access to advanced graphic design capabilities. However, it also raises concerns about the authenticity of visual content, posing significant challenges for image forensics (Ferreira et al., 2020), misinformation combating (Xu et al., 2023a), and copyright protection (Ren et al., 2024). In this paper, we consider the problem of distinguishing between images generated by AI models and those originating from real-world sources.
@@
In the literature, although there are numerous AI-generated image detectors (Wang et al., 2020; Frank et al., 2020; Ojha et al., 2023; Wang et al., 2023; Zhong et al., 2023; Ricker et al., 2024) and benchmarks (Wang et al., 2020; 2023; Zhu et al., 2024; Hong & Zhang, 2024), the prevailing problem formulation typically involves training models on images generated solely by GANs (e.g., ProGAN (Karras et al., 2017)) and evaluating their performance on datasets including images from various generative models, including GANs and DMs. However, such formulation poses two fundamental issues in practice. Firstly, evaluation benchmarks are simple, as they often feature test sets composed of random images from generative models, rather than images that present genuine challenges for human perception. Secondly, confining models to train exclusively on images from certain type of generative models (GANs or DMs) imposes an unrealistic constraint, hindering the model’s ability to learn from the diverse properties exhibited by more advanced generative models.
@@
To address the aforementioned issues, we propose two pivotal strategies. Firstly, we introduce a novel testset for AI-generated image detection, named Chameleon, manually annotated to include images that genuinely challenge human perception. This dataset has three key features: (i) Deceptively real: all AI-generated images in the dataset have passed a human perception ”Turing Test”, i.e., human annotators have misclassified them as real images. (ii) Diverse categories: comprising images of human, animal, object, and scene categories, the dataset depicts real-world scenarios across various contexts. (iii) High resolution: with most images having resolutions exceeding 720P and going up to 4K, all images in the dataset exhibit exceptional clarity. Consequently, this test set offers a more realistic evaluation of model performance. After evaluating 9 off-the-shelf AI-generated image detectors on Chameleon, unfortunately, all detectors suffer from significant performance drops, misclassifying the AI-generated images as real ones. Secondly, we reformulate the AI-generated image detection problem setup, which enables models to train across a broader spectrum of generative models, enhancing their adaptability and robustness in real-world scenarios.
@@
Based on the above observation, it is clear that detecting AI-generated images remains challenging, and is far from being solved. Therefore, a fundamental question arises: what distinguishes AI-generated images from real ones? Intuitively, such cues may appear from various aspects, including low-level textures or pixel statistics (e.g., the presence of white noise during image capturing), and high-level semantics (e.g., penguins are unlikely to be appearing on the grassland in Africa), geometry principle (e.g., perspective), physics (e.g., lighting condition). To reflect such intuition, we propose a simple AI-generated image detector, termed as AIDE (AI-generated Image DEtector with Hybrid Features). Specifically, AIDE incorporates a DCT (Ahmed et al., 1974) scoring module to capture low-level pixel statistics by extracting both high and low-frequency patches from the image, which are then processed through SRM (Spatial Rich Model) filters (Fridrich & Kodovsky, 2012) to characterize the noise pattern. Additionally, to capture global semantics, we utilize the pre-trained OpenCLIP (Ilharco et al., 2021) to encode the entire image. The features from various levels are fused in the channel dimension for the final prediction. To evaluate the effectiveness of our model, we conduct extensive experiments on two popular benchmarks, including AIGCDetectBenchmark (Wang et al., 2020) and GenImage (Zhu et al., 2024), for AI-generated image detection. On AIGCDetectBenchmark and GenImage benchmarks, AIDE surpasses state-of-the-art (SOTA) methods by +3.5% and +4.6% in accuracy scores, respectively. Moreover, AIDE also achieves competitive performance on our Chameleon benchmark.
@@
Overall, our contributions are summarized as follows: (i) We present the Chameleon dataset, a meticulously curated test set designed to challenge human perception by including images that deceptively resemble real-world scenes. With thorough evaluation of 9 different off-the-shelf detectors, this dataset exposes the limitations of existing approaches. (ii) We present a simple mixture-of-expert model, termed as AIDE, that enables to discern AI-generated images based on both low-level pixel statistics and high-level semantics. (iii) Experimentally, our model achieves state-of-the-art results on public benchmarks for AIGCDetectBenchmark (Wang et al., 2020) and GenImage (Zhu et al., 2024). While on Chameleon, it acts as a competitive baseline on a realistic evaluation benchmark, to foster future research in this community.
@@ 2 Related Works
AI-generated image detection. The demand for detecting AI-generated images has long been present. Early studies primarily focus on spatial domain cues, such as color (McCloskey & Albright, 2018), saturation (McCloskey & Albright, 2019), co-occurrence (Nataraj et al., 2019), and reflections (O’brien & Farid, 2012). However, these methods often suffer from limited generalization capabilities as generators progress. To address this limitation, CNNSpot (Wang et al., 2020) discovers that an image classifier trained exclusively on ProGAN (Karras et al., 2017) generator could generalize effectively to other unseen GAN architectures, with careful pre- and post-processing and data augmentation. FreDect (Frank et al., 2020) observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures. More recent approaches have explored novel perspectives for superior generalization ability. UnivFD (Ojha et al., 2023) proposes to train a universal liner classifier with pretrained CLIP-ViT (Dosovitskiy et al., 2020; Radford et al., 2021) features. DIRE (Wang et al., 2023) introduces DIRE features, which computes the difference between images and their reconstructions from pretrained ADM (Dhariwal & Nichol, 2021), to train a deep classifier. PatchCraft (Zhong et al., 2023) compares rich-texture and poor-texture patches from images, extracting the inter-pixel correlation discrepancy as a universal fingerprint, which is reported to achieve the state-of-the-art (SOTA) generalization performance. AEROBLADE (Ricker et al., 2024) proposes a training-free detection method for latent diffusion models using autoencoder reconstruction errors. However, these methods only discriminate real or fake images from a single perspective, often failing to generalize across images from different generators.
@@
AI-generated image datasets. To facilitate AI-generated image detection, many datasets containing both real and fake images have been organized for training and evaluation. Early dataset from CNNSpot (Wang et al., 2020) collects fake images from GAN series generators (Goodfellow et al., 2014; Zhu et al., 2017; Brock et al., 2018; Karras et al., 2019). Particularly, this dataset generates fake images exclusively using ProGAN (Karras et al., 2017) as training data and evaluates the generalization ability on a set of GAN-based testing data. However, with recent emergence of more advanced generators, such as diffusion model (DM) (Ho et al., 2020) and its variants (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Rombach et al., 2022; Song et al., 2020; Liu et al., 2022b; Lu et al., 2022; Hertz et al., 2022; Nichol et al., 2021), their realistic generations make visual differences between real and fake images progressively harder to detect. Subsequently, more datasets including DM-generated images have been proposed one after another, including DE-FAKE (Xu et al., 2023b), CiFAKE (Bird & Lotfi, 2024), DiffusionDB (Wang et al., 2022), ArtiFact (Rahman et al., 2023). One representative dataset is GenImage (Zhu et al., 2024), which comprises ImageNet’s 1,000 classes generated using 8 SOTA generators in both academia (e.g., Stable Diffusion (Sta, 2022)) and industry (e.g., Midjourney (mid, )). More recently, Hong et al. introduce a more comprehensive dataset, WildFake (Hong & Zhang, 2024), which includes AI-generated images sourced from multiple generators, architectures, weights, and versions. However, existing benchmarks only evaluate AI-generated images using current foundational models with simple prompts and few modifications, whereas deceptively real images from online communities usually necessitate hundreds to thousands of manual parameter adjustments.
@@ 3.1 Problem Formulation
In this paper, our goal is to train a computational model that can distinguish the AI-generated images from the ones captured by the camera, i.e., $y=\Phi_{\text{model}}(\mathbf{I};\Theta)\in\{0,1\}$, where $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ denotes an input RGB image, $\Theta$ refers to the learnable parameters. For training and testing, we consider the following two settings:
Train-Test Setting-I. In the literature, existing works on detecting AI-generated images (Wang et al., 2020; Frank et al., 2020; Ojha et al., 2023; Wang et al., 2023; Zhong et al., 2023) have exclusively considered the scenario of training on images from single generative model, for example, ProGAN (Karras et al., 2017), or Stable Diffusion (Sta, 2022), and then evaluated on images from various generative models. That is,
$\mathcal{G}_{\text{train}}=\mathcal{G}_{\text{GAN}}\vee\mathcal{G}_{\text{DM}},\mathcal{G}_{\text{test}}=\{\mathcal{G}_{\text{ProGAN}},\mathcal{G}_{\text{CycleGAN}},...,\mathcal{G}_{\text{SD}},\mathcal{G}_{\text{Midjourney}}\}.$ (1)
@@
$\mathcal{G}_{\text{train}}=\{\mathcal{G}_{\text{GAN}},\mathcal{G}_{\text{DM}}\},\mathcal{G}_{\text{test}}=\{\mathcal{D}_{\texttt{Chameleon}}\}.$ (2)
$\mathcal{D}_{\texttt{Chameleon}}$ refers to our proposed benchmark, as detailed below. We believe this setting resembles more practical scenario for future model development in this community.
@@ 3.2 Chameleon Dataset
The primary objective of the Chameleon dataset is to evaluate the generalization and robustness of existing AI-generated image detectors, for a sanity check on the progress of AI-generated image detection. In this section, we outline the progression of the proposed dataset in three critical phases: (i) dataset collection, (ii) dataset curation, (iii) dataset annotation, and (iv) dataset comparison. The statistical results of our dataset are illustrated in Table 1 and we compare our dataset with existing benchmarks in Fig. 1.
@@
Real Images Fake Images
Scene 3,574 2,976
Object 3,578 2,016
Animal 3,998 313
Human 3,713 5,865
Total 14,863 11,170
@@ 3.2.1 Dataset Collection
To simulate real-world cases on detecting AI-generated images, we structure our Chameleon dataset based on three main principles: (i) images must be deceptively real, and (ii) they should cover a diverse range of categories, and (iii) they should also have very high image quality. Importantly, each image must have (a) Creative Commons (CC BY 4.0) license, or (b) explicit permissions obtained from the owners to use in our research. Herein, we present the details of image collection.
Fake Image Collection: To collect images that are deceptively real, and cover sufficiently diverse categories, we source user-created AI-generated images from popular AI-painting communities (i.e., ArtStation (art, ), Civitai (civ, ), and Liblib (lib, )), many of which utilize commercial APIs (e.g., Midjourney (mid, ) and DALLE-3 (Ramesh et al., 2022)) or various LoRA modules (Hu et al., 2021) with Stable Diffusion (SD) (Sta, 2022) that fine-tuned on their in-house data. Specifically, we initiate the process by utilizing GPT-4 (cha, 2022) to generate diverse query words to retrieve AI-generated images. Throughout the collection process, we enforce stringent NSFW (Not Safe For Work) restrictions. Ultimately, our collection comprises over 150K fake images.
Real Image Collection: To ensure that real and fake images fall into the same distribution, we employ identical query words to retrieve real images, mirroring the approach used for gathering AI-generated images. Eventually, we collect over 20K images from platforms like Unsplash (uns, ), which is an online community providing high-quality, free-to-use images contributed by photographers worldwide.

### v3

@@ Abstract
With the rapid development of generative models, discerning AI-generated content has evoked increasing attention from both industry and academia. In this paper, we conduct a sanity check on whether the task of AI-generated image detection has been solved. To start with, we present Chameleon dataset, consisting of AI-generated images that are genuinely challenging for human perception. To quantify the generalization of existing methods, we evaluate 9 off-the-shelf AI-generated image detectors on Chameleon dataset. Upon analysis, almost all models misclassify AI-generated images as real ones. Later, we propose AIDE (AI-generated Image DEtector with Hybrid Features), which leverages multiple experts to simultaneously extract visual artifacts and noise patterns. Specifically, to capture the high-level semantics, we utilize CLIP to compute the visual embedding. This effectively enables the model to discern AI-generated images based on semantics and contextual information. Secondly, we select the highest and lowest frequency patches in the image, and compute the low-level patchwise features, aiming to detect AI-generated images by low-level artifacts, for example, noise patterns, anti-aliasing effects. While evaluating on existing benchmarks, for example, AIGCDetectBenchmark and GenImage, AIDE achieves +3.5% and +4.6% improvements to state-of-the-art methods, and on our proposed challenging Chameleon benchmarks, it also achieves promising results, despite the problem of detecting AI-generated images remains far from being solved.
* Equal contribution. $\dagger$ Corresponding author.
@@ 1 Introduction
Recently, the vision community has witnessed remarkable advancements in generative models. These methods, ranging from generative adversarial networks (GANs) (Goodfellow et al., 2014; Zhu et al., 2017; Brock et al., 2019; Karras et al., 2019) to diffusion models (DMs) (Ho et al., 2020; Nichol & Dhariwal, 2021; Rombach et al., 2022; Song et al., 2021; Liu et al., 2022b; Lu et al., 2022; Hertz et al., 2023; Nichol et al., 2022) have demonstrated unprecedented capabilities in synthesizing high-quality images that closely resemble real-world scenes. On the positive side, such generative models have enabled various valuable tools for artists and designers, democratizing access to advanced graphic design capabilities. However, it also raises concerns about the authenticity of visual content, posing significant challenges for image forensics (Ferreira et al., 2020), misinformation combating (Xu et al., 2023a), and copyright protection (Ren et al., 2024). In this paper, we consider the problem of distinguishing between images generated by AI models and those originating from real-world sources.
@@
In the literature, although there are numerous AI-generated image detectors (Wang et al., 2020; Frank et al., 2020; Ojha et al., 2023; Wang et al., 2023a; Zhong et al., 2023; Ricker et al., 2024) and benchmarks (Wang et al., 2020; 2023a; Zhu et al., 2024; Hong & Zhang, 2024), the prevailing problem formulation typically involves training models on images generated solely by GANs (e.g., ProGAN (Karras et al., 2018)) and evaluating their performance on datasets including images from various generative models, including GANs and DMs. However, such formulation poses two fundamental issues in practice: (i) evaluation benchmarks are simple, as they often feature test sets composed of random images from generative models, rather than images that present genuine challenges for human perception; (ii) confining models to train exclusively on images from certain types of generative models (GANs or DMs) imposes an unrealistic constraint, hindering the model’s ability to learn from the diverse properties exhibited by more advanced generative models.
@@
To address the aforementioned issues, we propose two pivotal strategies. Firstly, we introduce a novel testset for AI-generated image detection, named Chameleon, manually annotated to include images that genuinely challenge human perception. This dataset has three key features: (i) deceptively real: all AI-generated images in the dataset have passed a human perception “Turing Test”, i.e., human annotators have misclassified them as real images. (ii) diverse categories: comprising images of human, animal, object, and scene categories, the dataset depicts real-world scenarios across various contexts. (iii) high resolution: with most images having resolutions exceeding 720P and going up to 4K, all images in the dataset exhibit exceptional clarity. Consequently, this test set offers a more realistic evaluation of model performance. After evaluating 9 off-the-shelf AI-generated image detectors on Chameleon, unfortunately, all detectors suffer from significant performance drops, mis-classifying the AI-generated images as real ones. Secondly, we reformulate the AI-generated image detection problem setup, which enables models to train across a broader spectrum of generative models, enhancing their adaptability and robustness in real-world scenarios.
@@
Based on the above observation, it is clear that detecting AI-generated images remains challenging, and far from being solved. Therefore, a fundamental question arises: what distinguishes AI-generated images from real ones? Intuitively, such cues may appear from various aspects, including low-level textures or pixel statistics (e.g., the presence of white noise during image capturing), and high-level semantics (e.g., penguins are unlikely to be appearing on the grassland in Africa), geometry principle (e.g., perspective), physics (e.g., lighting condition). To reflect such intuition, we propose a simple AI-generated image detector, termed as AIDE (AI-generated Image DEtector with Hybrid Features). Specifically, AIDE incorporates a DCT (Ahmed et al., 1974) scoring module to capture low-level pixel statistics by extracting both high and low-frequency patches from the image, which are then processed through SRM (Spatial Rich Model) filters (Fridrich & Kodovsky, 2012) to characterize the noise pattern. Additionally, to capture global semantics, we utilize the pre-trained OpenCLIP (Ilharco et al., 2021) to encode the entire image. The features from various levels are fused in the channel dimension for the final prediction. To evaluate the effectiveness of our model, we conduct extensive experiments on two popular benchmarks, including AIGCDetectBenchmark (Wang et al., 2020) and GenImage (Zhu et al., 2024), for AI-generated image detection. On AIGCDetectBenchmark and GenImage benchmarks, AIDE surpasses state-of-the-art (SOTA) methods by +3.5% and +4.6% in accuracy scores, respectively. Moreover, AIDE also achieves competitive performance on our Chameleon benchmark.
@@
Overall, our contributions are summarized as follows: (i) we present the Chameleon dataset, a meticulously curated test set designed to challenge human perception by including images that deceptively resemble real-world scenes. With thorough evaluation of 9 different off-the-shelf detectors, this dataset exposes the limitations of existing approaches; (ii) we present a simple mixture-of-expert model, termed AIDE, that enables to discern AI-generated images based on both low-level pixel statistics and high-level semantics; (iii) experimentally, our model achieves state-of-the-art results on public benchmarks for AIGCDetectBenchmark (Wang et al., 2020) and GenImage (Zhu et al., 2024). While on Chameleon, it acts as a competitive baseline on a realistic evaluation benchmark, to foster future research in this community.
@@ 2 Related Works
AI-generated Image Detection. The demand for detecting AI-generated images has long been present. Early studies primarily focus on spatial domain cues, such as color (McCloskey & Albright, 2018), saturation (McCloskey & Albright, 2019), co-occurrence (Nataraj et al., 2019), and reflections (O’brien & Farid, 2012). However, these methods often suffer from limited generalization capabilities as generators progress. To address this limitation, CNNSpot (Wang et al., 2020) discovers that an image classifier trained exclusively on ProGAN (Karras et al., 2018) generator could generalize effectively to other unseen GAN architectures, with careful pre- and post-processing and data augmentation. FreDect (Frank et al., 2020) observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures. More recent approaches have explored novel perspectives for superior generalization ability. UnivFD (Ojha et al., 2023) proposes to train a universal liner classifier with pretrained CLIP-ViT (Dosovitskiy et al., 2020; Radford et al., 2021) features. DIRE (Wang et al., 2023a) introduces DIRE features, which computes the difference between images and their reconstructions from pretrained ADM (Dhariwal & Nichol, 2021), to train a deep classifier. PatchCraft (Zhong et al., 2023) compares rich-texture and poor-texture patches from images, extracting the inter-pixel correlation discrepancy as a universal fingerprint, which is reported to achieve the state-of-the-art (SOTA) generalization performance. AEROBLADE (Ricker et al., 2024) proposes a training-free detection method for latent diffusion models using autoencoder reconstruction errors. FatFormer (Liu et al., 2024a) introduces a forgery-aware adapter to discern and integrate local forgery traces based on CLIP. CLIPMoLE (Liu et al., 2024b) adapts a combination of shared and separate LoRAs within an MoE-based structure in deeper ViT blocks. SSP (Chen et al., 2024) employs the simplest patches for detection. However, these methods only discriminate real or fake images from a single perspective, often failing to generalize across images from different generators.
@@
AI-generated Image Datasets. To facilitate AI-generated image detection, many datasets containing both real and fake images have been organized for training and evaluation. Early dataset from CNNSpot (Wang et al., 2020) collects fake images from GAN series generators (Goodfellow et al., 2014; Zhu et al., 2017; Brock et al., 2019; Karras et al., 2019). Particularly, this dataset generates fake images exclusively using ProGAN (Karras et al., 2018) as training data and evaluates the generalization ability on a set of GAN-based testing data. However, with recent emergence of more advanced generators, such as diffusion model (DM) (Ho et al., 2020) and its variants (Dhariwal & Nichol, 2021; Nichol & Dhariwal, 2021; Rombach et al., 2022; Song et al., 2021; Liu et al., 2022b; Lu et al., 2022; Hertz et al., 2023; Nichol et al., 2022), their realistic generations make visual differences between real and fake images progressively harder to detect. Subsequently, more datasets including DM-generated images have been proposed one after another, including DE-FAKE (Xu et al., 2023b), CiFAKE (Bird & Lotfi, 2024), DiffusionDB (Wang et al., 2023b), ArtiFact (Rahman et al., 2023). One representative dataset is GenImage (Zhu et al., 2024), which comprises ImageNet’s 1,000 classes generated using 8 SOTA generators in both academia (e.g., Stable Diffusion (Sta, 2022)) and industry (e.g., Midjourney (mid, )). More recently, Hong et al. introduce a more comprehensive dataset, WildFake (Hong & Zhang, 2024), which includes AI-generated images sourced from multiple generators, architectures, weights, and versions. However, existing benchmarks only evaluate AI-generated images using current foundational models with simple prompts and few modifications, whereas deceptively real images from online communities usually necessitate hundreds to thousands of manual parameter adjustments.
@@ 3.1 Problem Formulation
In this paper, our goal is to train a model that can distinguish the AI-generated images from the ones captured by the camera, i.e., $y=\Phi_{\text{model}}(\mathbf{I};\Theta)\in\{0,1\}$, where $\mathbf{I}\in\mathbb{R}^{H\times W\times 3}$ denotes an input RGB image, $\Theta$ refers to the learnable parameters. For training and testing, we consider the following two settings:
Train-Test Setting-I. In the literature, existing works on detecting AI-generated images (Wang et al., 2020; Frank et al., 2020; Ojha et al., 2023; Wang et al., 2023a; Zhong et al., 2023) have exclusively considered the scenario of training on images from single generative model, for example, ProGAN (Karras et al., 2018), or Stable Diffusion (Sta, 2022), and then evaluated on images from various generative models. That is,
$\mathcal{G}_{\text{train}}=\mathcal{G}_{\text{GAN}}\vee\mathcal{G}_{\text{DM}},\quad\mathcal{G}_{\text{test}}=\{\mathcal{G}_{\text{ProGAN}},\mathcal{G}_{\text{CycleGAN}},...,\mathcal{G}_{\text{SD}},\mathcal{G}_{\text{Midjourney}}\}$ (1)
@@
$\mathcal{G}_{\text{train}}=\{\mathcal{G}_{\text{GAN}},\mathcal{G}_{\text{DM}}\},\quad\mathcal{G}_{\text{test}}=\{\mathcal{D}_{\texttt{Chameleon}}\}$ (2)
$\mathcal{D}_{\texttt{Chameleon}}$ refers to our proposed benchmark, as detailed below. We believe this setting resembles a more practical scenario for future model development in this community.
@@ 3.2 Chameleon Dataset
The primary objective of the Chameleon dataset is to evaluate the generalization and robustness of existing AI-generated image detectors, for a sanity check on the progress of AI-generated image detection. In this section, we outline the progression of the proposed dataset, including: (i) collection, (ii) curation, (iii) annotation. The statistical results of our dataset are illustrated in Table 1 and we compare our dataset with existing benchmarks in Fig. 1.
@@
Scene Object Animal Human Total
Real Images 3,574 3,578 3,998 3,713 14,863
Fake Images 2,976 2,016 313 5,865 11,170
@@ 3.2.1 Dataset Collection
To simulate the real-world scenarios in detecting AI-generated images, we structure Chameleon dataset based on three main principles: (i) images must be deceptively real, and (ii) they should cover a diverse range of categories, and (iii) they should also have very high image quality. Importantly, each image must have (a) Creative Commons (CC BY 4.0) license, or (b) explicit permissions obtained from the owners to use in our research. Herein, we present the details of image collection.
Fake Image Collection: To collect images that are deceptively real, and cover sufficiently diverse categories, we source user-created AI-generated images from popular AI-painting communities, i.e., ArtStation (art, ), Civitai (civ, ), and Liblib (lib, ), many of which utilize commercial APIs (e.g., Midjourney (mid, ) and DALLE-3 (Ramesh et al., 2022)) or various LoRA modules (Hu et al., 2022) with Stable Diffusion (SD) (Sta, 2022) that fine-tuned on their in-house data. Specifically, we initiate the process by utilizing GPT-4 (cha, 2022) to generate diverse query words to retrieve AI-generated images. Throughout the collection process, we enforce stringent NSFW (Not Safe For Work) restrictions. Ultimately, our collection comprises over 150K AI-generated images.
Real Image Collection: To ensure that real and AI-generated images fall into the same distribution, we employ identical query words to retrieve real images, mirroring the approach used for gathering AI-generated images. Eventually, we collect over 20K images from platforms like Unsplash (uns, ), which is an online community providing high-quality, free-to-use images contributed by photographers worldwide.